OpenStudy (anonymous):

The acceleration of a particle is a(t) = 4 cos (2t). Find the position function for the particle if v(0) = -1 and s(0) = -3

OpenStudy (anonymous):

Integrate the function and use the points to find the constant of integration

OpenStudy (anonymous):

like @PhysicsGuru said \[a(t)=\frac{\mathrm{d}v(t)}{\mathrm{d}t}\] \[\int \mathrm{d}v= \int a(t) \ \mathrm{d}t \] \[v(t)+v_0=\int a(t) \ \mathrm{d}t\] \[v(t)=\frac{\mathrm{d}s(t)}{\mathrm{d}t}\] \[ \int \mathrm{d}s(t) = \int \big(v(t)+v_0\big) \mathrm{d}t\] also \[ \int \cos (2t) \mathrm{d} t = \frac{1}{2} \sin (2t) +C\] \[ \int \sin(2t) = -\frac{1}{2} \cos(2t)+C\]